OpenStudy (anonymous):

Jonah is going to the store to buy candles. Small candles cost $3.50 and large candles cost $5. He needs to buy at least 20 candles, and he cannot spend more than $80. Write a system of linear inequalities that represent the situation.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

3.50x+5y<=80 y>=80?

jimthompson5910 (jim_thompson5910):

`3.50x+5y<=80` looks good

OpenStudy (anonymous):

Ok

jimthompson5910 (jim_thompson5910):

x = number of small candles y = number of large candles x+y = total number of candles x+y >= 20 because `He needs to buy at least 20 candles`

OpenStudy (anonymous):

Thank you! I knew I did something wrong. Thanks for finding my mistake.

jimthompson5910 (jim_thompson5910):

so the system is \[\Large \begin{cases}3.5x+5y\le80\\x+y\ge 20\\x\ge0\\y\ge0\end{cases}\] x>=0 and y >=0 are thrown in there to force x and y to be nonnegative

OpenStudy (anonymous):

Ok thank you!